# Modality-Aware Neuron Pruning for Unlearning in Multimodal Large Language Models

### v1

@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023); Fu et al. (2024); Qin et al. (2023) and Multimodal Large Language Models (MLLMs) Liu et al. (2024a); Ye et al. (2023, 2024); Zhu et al. (2023); Zhang et al. (2025) have showcased their exceptional capabilities across various AI domains Ouyang et al. (2022); Tan et al. (2024); Ni et al. (2025); Zhang et al. (2024b), largely due to extensive pre-training and fine-tuning on vast data corpus. However, this remarkable learning ability also poses risks such as privacy violations and copyright infringements. Since retraining from scratch while excluding these data is computationally expensive, Machine Unlearning (MU) Nguyen et al. (2022); Liu et al. (2024c, f) has emerged as an efficient alternative to remove the influence of sensitive data while preserving overall model performance.
@@ 5.2 Pruning Ratio Analysis
From the figures, we observe that MANU maintains a robust balance between unlearning performance and model utility across various aspects. The better an algorithm balances these two aspects, the closer it will appear to the top-right in the figure, indicating a larger difference in forget accuracy and higher retain accuracy. For example, in Figures 5(a) and 5(b), MANU achieves a comparable reduction in Forget Set accuracy to GA-based approaches while maintaining high accuracy on the Retain and Real Celebrity sets. Similarly, when evaluating model reasoning abilities using MMMU and Llava-Bench (i.e. Figures 5(c) and 5(d)), MANU performs comparably to prompting techniques while significantly surpassing them in forget set accuracy. Thus, MANU effectively balances unlearning and model utility across multiple dimensions. Further analysis can be found in Appendix E.4.
@@ Appendix C Rationale for Targeting MLP Layers
Recent research has demonstrated that MLP layers serve as primary knowledge storage components in transformer architectures. For example, Huang et al. (2024) introduces the concept of "knowledge neurons," highlighting that specific neurons within MLP layers are responsible for encoding and storing information. By manipulating these neurons, it is possible to edit or selectively remove knowledge, offering fine-grained control over the model’s retained information. Beyond individual neurons, broader findings in knowledge editing literature reinforce the significance of MLP layers for model knowledge control. Prior works have shown that knowledge manipulation techniques, including direct parameter modification and knowledge attribution methods, consistently identify MLP layers as the primary repository of factual and task-specific knowledge Wang et al. (2024); Meng et al. (2022a, b). Given that vision transformers share a fundamentally similar architecture with language transformers, where MLPs play an analogous role in feature extraction and information processing Ghiasi et al. (2022), we extend this insight to the vision tower as well. The effectiveness of targeting MLPs for knowledge controlling is further supported by recent work in LLM pruning Pochinkov and Schoots (2024), which demonstrates that modifying MLP layers enables precise control over model knowledge while maintaining core model capabilities.

### v2

@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023); Fu et al. (2024); Qin et al. (2023) and Multimodal Large Language Models (MLLMs) Liu et al. (2024a); Ye et al. (2023, 2024); Zhu et al. (2023); Zhang et al. (2025b, a) have showcased their exceptional capabilities across various AI domains Ouyang et al. (2022); Tan et al. (2024); Ni et al. (2025); Zhang et al. (2024b), largely due to extensive pre-training and fine-tuning on vast data corpus. However, this remarkable learning ability also poses risks such as privacy violations and copyright infringements. Since retraining from scratch while excluding these data is computationally expensive, Machine Unlearning (MU) Nguyen et al. (2022); Wang et al. (2024b); Liu et al. (2024c, f) has emerged as an efficient alternative to remove the influence of sensitive data while preserving overall model performance.
@@ 5.2 Pruning Ratio Analysis
From the figures, we observe that MANU maintains a robust balance between unlearning performance and model utility across various aspects. The better an algorithm balances these two aspects, the closer it will appear to the top-right in the figure, indicating a larger difference in forget accuracy and higher retain accuracy. For example, in Figures 5(a) and 5(b), MANU achieves a comparable reduction in Forget Set accuracy to GA-based approaches while maintaining high accuracy on the Retain and Real Celebrity sets. Similarly, when evaluating model reasoning abilities using MMMU and Llava-Bench (i.e. Figures 5(c) and 5(d)), MANU performs comparably to prompting techniques while significantly surpassing them in forget set accuracy. Thus, MANU effectively balances unlearning and model utility across multiple dimensions. Further analysis and additional ablation studies can be found in Appendix E.4 and E.6.
@@ Appendix C Rationale for Targeting MLP Layers
Recent research has demonstrated that MLP layers serve as primary knowledge storage components in transformer architectures. For example, Huang et al. (2024) introduces the concept of "knowledge neurons," highlighting that specific neurons within MLP layers are responsible for encoding and storing information. By manipulating these neurons, it is possible to edit or selectively remove knowledge, offering fine-grained control over the model’s retained information. Beyond individual neurons, broader findings in knowledge editing literature reinforce the significance of MLP layers for model knowledge control. Prior works have shown that knowledge manipulation techniques, including direct parameter modification and knowledge attribution methods, consistently identify MLP layers as the primary repository of factual and task-specific knowledge Wang et al. (2024a); Meng et al. (2022a, b). Given that vision transformers share a fundamentally similar architecture with language transformers, where MLPs play an analogous role in feature extraction and information processing Ghiasi et al. (2022), we extend this insight to the vision tower as well. The effectiveness of targeting MLPs for knowledge controlling is further supported by recent work in LLM pruning Pochinkov and Schoots (2024), which demonstrates that modifying MLP layers enables precise control over model knowledge while maintaining core model capabilities.
@@ E.3 Unlearning across modalities
E.5 Appendix: Additional Utility Datasets
In addition to the evaluations already included—specifically on MMMU and LLaVA-Bench, we present results on three additional downstream benchmarks to further assess the functional utility of MANU across a diverse range of multimodal tasks. These benchmarks include: MIA-Bench Qian et al. (2024) for evaluating conversational abilities, MM-Vet Yu et al. (2023) for assessing integrated multimodal reasoning, and MMStar Chen et al. (2024) for testing vision-indispensable capabilities.
In Table 6, we report the average performance of the vanilla model and unlearned models using various unlearning methods on these benchmarks. As shown in the Table, MANU performs competitively across all three benchmarks, often outperforming or matching the baselines. These results suggest that MANU effectively preserves cross-modal alignment and functional utility, even after selective neuron pruning.
LLaVA (5% Forget Set)
Model MIA-Bench (Avg) MM-Vet (Avg) MMStar (Avg)
Vanilla 69.60 34.08 34.80
GA 57.29 28.92 29.83
Grad. Diff 59.02 30.28 30.13
KL Minim. 62.24 31.29 31.29
NPO 63.44 32.38 32.32
MANU 64.02 32.76 32.41
Idefics2 (5% Forget Set)
Model MIA-Bench (Avg) MM-Vet (Avg) MMStar (Avg)
Vanilla 54.42 43.90 49.27
GA 47.77 35.12 40.90
Grad. Diff 48.58 36.67 41.02
KL Minim. 49.64 38.61 41.09
NPO 50.89 40.22 42.88
MANU 51.33 40.83 44.71
Table 6: Evaluation results on MIA-Bench, MM-Vet, and MMStar for LLaVA and Idefics2 under the 5% Forget Set setting. Higher scores indicate better performance.
E.6 Appendix: Ablations on Importance Functions
To further investigate the contribution of each importance function, here we provide additional ablation studies where we iteratively zeroed out each importance function in the scoring formula. The results are displayed in Table 7. From the ablation results, we observe that each importance function contributes uniquely to the overall effectiveness of MANU, and removing any single component results in a noticeable trade-off between unlearning performance and utility preservation. Since the trends for LLaVA and Idefics2 are consistent, we use the LLaVA results as a representative example.
Specifically, removing Frequency Importance ($I_{\text{freq}}$) or Variance Importance ($I_{\text{var}}$) substantially worsens unlearning on the Forget and Test sets—e.g., classification accuracy rises from 41.25% (MANU) to 47.35% and 46.67%, respectively, indicating a failure to sufficiently erase the target knowledge. These two metrics are particularly valuable for identifying neurons consistently and distinctively activated by forget data, thus supporting targeted unlearning. On the other hand, removing Absolute Importance ($I_{\text{abs}}$) or RMS Importance ($I_{\text{rms}}$) more prominently degrades performance on the Retain and Real Celebrity sets. For instance, when $I_{\text{abs}}$ is excluded, Retain classification accuracy drops from 43.38% to 42.37%, and Real Celebrity classification accuracy declines to 46.59%. This suggests that $I_{\text{abs}}$ and $I_{\text{rms}}$ are important for preserving high-activation neurons that contribute broadly to general reasoning, thus maintaining utility. These findings support our equal-weighting strategy, where each importance score captures a distinct and complementary signal. While we acknowledge that learned or tuned weightings might yield further improvements, we leave such model-specific enhancements for future work.
Models Forget Set Test Set Retain Set Real Celebrity
Class.
Acc ($\downarrow$)
Rouge
Score ($\downarrow$)
Fact.
Score ($\downarrow$)
Cloze
Acc ($\downarrow$)
Class.
Acc ($\downarrow$)
Rouge
Score ($\downarrow$)
Fact.
Score ($\downarrow$)
Cloze
Acc ($\downarrow$)
Class.
Acc ($\uparrow$)
Rouge
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Cloze
Acc ($\uparrow$)
Class.
Acc ($\uparrow$)
Rouge
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Cloze
Acc ($\uparrow$)
LLaVA-7B (5% Forget)
Vanilla 51.70 0.645 6.78 25.81 47.86 0.539 4.89 23.01 46.11 0.632 6.41 27.83 51.80 0.479 5.47 17.35
w/o $I_{abs}$ 40.33 0.472 3.30 16.83 41.67 0.354 3.62 14.15 42.37 0.458 4.23 19.22 46.59 0.375 4.12 13.38
w/o $I_{freq}$ 47.35 0.612 4.09 25.01 45.67 0.472 4.22 21.78 45.12 0.522 5.16 24.16 49.86 0.459 4.99 15.73
w/o $I_{var}$ 46.67 0.636 4.14 24.92 44.17 0.460 4.04 22.75 44.09 0.564 5.39 24.64 50.61 0.461 4.85 16.07
w/o $I_{rms}$ 40.82 0.500 3.19 15.17 40.33 0.382 3.70 16.03 41.02 0.471 4.19 20.68 46.08 0.466 4.05 12.03
MANU 41.25 0.491 3.27 17.08 41.67 0.334 3.81 15.78 43.38 0.542 4.45 24.08 49.57 0.448 4.67 16.01
Idefics2-8B (5% Forget)
Vanilla 53.80 0.630 6.22 4.75 47.86 0.434 5.00 24.97 46.11 0.644 6.51 42.35 52.75 0.459 5.75 20.05
w/o $I_{abs}$ 35.98 0.403 3.17 3.32 33.86 0.354 3.89 18.55 41.06 0.542 4.50 39.22 48.82 0.366 3.92 14.91
w/o $I_{freq}$ 48.81 0.608 4.72 4.41 42.39 0.407 4.07 23.49 44.14 0.583 5.87 41.09 51.54 0.428 5.09 19.33
w/o $I_{var}$ 46.65 0.588 4.66 4.83 44.77 0.395 4.14 23.82 45.34 0.626 5.80 41.43 52.22 0.417 4.99 19.15
w/o $I_{rms}$ 36.18 0.420 3.05 2.18 32.19 0.324 3.32 17.77 40.91 0.507 4.44 37.19 46.88 0.342 3.73 14.11
MANU (full) 37.13 0.413 3.20 2.11 35.55 0.361 3.91 20.97 42.71 0.538 4.60 40.01 50.09 0.399 4.11 18.80
Table 7: Ablation study of MANU on two base MLLM models under a 5% forget data setup. Lower scores on the Forget/Test sets indicate better unlearning, while higher scores on the Retain/Celebrity sets indicate better utility preservation.
E.7 Appendix: Generalizability with Larger MLLMs
Model Forget Set Test Set Retain Set Real Celebrity
Class.
Acc ($\downarrow$)
Rouge
Score ($\downarrow$)
Fact.
Score ($\downarrow$)
Cloze
Acc ($\downarrow$)
Class.
Acc ($\downarrow$)
Rouge
Score ($\downarrow$)
Fact.
Score ($\downarrow$)
Cloze
Acc ($\downarrow$)
Class.
Acc ($\uparrow$)
Rouge
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Cloze
Acc ($\uparrow$)
Class.
Acc ($\uparrow$)
Rouge
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Cloze
Acc ($\uparrow$)
LLaVA-13B (5% Forget Set)
Vanilla 44.40 0.604 6.10 21.90 36.80 0.493 4.72 20.82 46.11 0.657 6.38 27.80 68.89 0.633 7.01 26.80
GA 34.04 0.423 2.98 15.19 26.35 0.386 3.74 15.01 35.69 0.477 3.04 16.94 48.84 0.447 4.56 18.02
Grad. Diff 34.16 0.435 3.95 15.55 25.99 0.362 3.80 16.13 37.54 0.498 3.55 18.80 49.25 0.493 4.97 18.34
KL Minim. 38.64 0.561 4.17 19.97 31.06 0.432 4.13 18.62 40.00 0.571 4.18 20.07 55.29 0.555 5.44 22.15
Prompting 42.99 0.588 5.36 20.80 34.80 0.454 4.52 20.91 42.81 0.622 5.23 27.66 67.72 0.633 6.93 25.76
NPO 35.12 0.473 3.41 18.77 29.73 0.397 3.89 17.77 40.79 0.565 4.55 20.23 59.83 0.579 5.86 22.54
MANU 33.65 0.434 3.13 15.33 25.41 0.382 3.70 16.08 42.88 0.593 4.72 22.75 59.97 0.585 6.08 22.80
Table 8: Overall results of MANU with varying pruning ratios on LLaVA-13B model under a 5% forget data setup. Lower scores on the Forget/Test sets indicate better unlearning, while higher scores on the Retain/Celebrity sets indicate better utility preservation.
To further evaluate the scalability and generalizability of MANU to larger MLLMs, we conducted an additional set of experiments using the LLaVA-13B architecture. In this section, we only report results for the 5% Forget Set split for reference, which required fine-tuning a separate vanilla LLaVA-13B model on the MLLMU-Bench dataset. The quantitative results are presented in Table 8.
As it shown in the table, the observed performance trends on LLaVA-13B are consistent with those reported for the smaller 7B and 8B variants. Specifically, MANU consistently ranks among the top-performing methods across all evaluation tasks. While GA and Gradient Difference occasionally achieve marginally better scores in terms of unlearning effectiveness, these methods generally underperform in preserving model utility. Conversely, the Prompting-based approach demonstrates strong utility preservation but exhibits significantly lower forgetting capability. MANU offers a robust compromise, maintaining competitive unlearning performance while preserving downstream utility across all evaluation settings.
These findings reaffirm the central design hypothesis behind MANU —that modality-specific importance signals can be effectively extracted and leveraged even within larger, more entangled model architectures. Additionally, we note that italicized emphasis used in our main tables to indicate second-best values may not be easily distinguishable in print. In the revised manuscript, we will replace italics with underlined formatting for improved visual clarity. This extension strengthens the empirical validation of MANU and demonstrates its applicability to state-of-the-art, large-scale MLLMs.